# Bootstrapping the critical behavior of multi-matrix models

### v1

@@ 3 The quartic model
Figure 1: Bootstrapped solution of the quartic Hermitian matrix model for $g>0$. The colours correspond to the the size of the submatrix of the Hankel matrix as follows: Light blue is for two by two, dark blue is for three by three, green is for four by four, and gold is for five by five. The black curve is the analytic solution.
@@
Figure 2: Bootstrapped solutions of the quartic Hermitian matrix model for $g<0$. The top left yellow region was computed with submatrices of size six by six, the top right was computed with submatrices of size seven by seven, and the bottom region was computed with submatrices of size ten by ten. The critical point of this model is $-\frac{1}{12}=-.0833\overline{3}$.
@@
The derivative of the free energy of the matrix model can then be expanded as
@@ 4 The 2-matrix models
This model can clearly be considered as a convergent integral when the coupling constant $g,\alpha,\beta\geq 0$, or as a formal matrix integral that enumerates coloured maps otherwise. As far as the authors are aware, there is no known technique that can be used to solve this model in its full generality, but it has been solved in special cases. When $(g,\alpha,\beta)=(g,0,0)$, this is two uncoupled quartic ensembles. When $(g,\alpha,\beta)=(g,\alpha,0)$, this becomes the symmetric $ABAB$ model solved via the method of characteristic expansion in [kazakov1999two]. When $(g,\alpha,\beta)=(0,0,\beta)$, this matrix integral is sometimes known as the Hoppe model and has been solved via a reduction to the KP equation or saddle point method [kazakov1999d, berenstein2009multi]. For the particular configuration $(g,\alpha,\beta)=(g,\alpha,\alpha)$, this model was bootstrapped in [kazakov2022analytic], where a phase diagram was constructed.
@@ 4.1 When $(g,\alpha,\beta)=(g,g,g)$
By symbolically solving SDE of the model in Mathematica, all the the mixed moments we examined could be expressed solely in terms of the coupling constant $g$ and $m_{2}$; see Appendix A.2 for some examples. This suggests that this model has a search space of one. Despite the uncertain nature of this claim, by using the solutions mentioned above, we are able to produce several analytical results and conjectures in addition to numerical bootstrap estimates.
@@ 4.1.1 Bootstrap bounds
The second moment $m_{2}$ of the matrix model 8 is such that
$\frac{1}{8g}-\frac{1}{8}\sqrt{\frac{16g+1}{g^{2}}}\leq m_{2}\leq\frac{1}{8g}+\frac{1}{8}\sqrt{\frac{16g+1}{g^{2}}},$
for $g\in[1/16,0)$. Additionally,
$-\frac{1}{8}\sqrt{\frac{16g+1}{g^{2}}}-\frac{1}{8g}<m_{2}<0,$
@@
$0<m_{2}<\frac{1}{8}\sqrt{\frac{16g+1}{g^{2}}}-\frac{1}{8g}$
@@
$\det\begin{bmatrix}1&0&0&m_{2}&0\\ 0&m_{2}&0&0&0\\ 0&0&m_{2}&0&0\\ m_{2}&0&0&m_{4}&0\\ 0&0&0&0&m_{ABAB}\end{bmatrix}\geq 0,$
@@
Figure 3: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (6) overlaid as follows: yellow is five by five, cyan is nine by nine, and purple is twenty-one by twenty-one. The subfigure on the right is only the latter constraint.
@@ 4.1.2 Series expansion at zero
$-\frac{d}{dg}F_{0}=\frac{1}{2}m_{4}+\frac{1}{2}m_{ABAB}+m_{AABB}.$
@@ 4.1.3 Critical behaviour
Figure 4: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model for $g<0$. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (6) as follows: yellow is five by five, cyan is nine by nine, and purple is twenty-one by twenty-one.
Figure 5: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model for $g<0$. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (6) as follows: yellow is five by five, cyan is nine by nine, and purple is twenty-one by twenty-one. This is compared with the estimates from the series expansion of the second moment in Section 4.1.2. Each coloured line from left to right represents successively more terms in expansion (9).
@@
Figure 6: Bootstrapped solution compared with equation (11), in red, for an estimate of $g_{c}=-0.0502729$. As in other plots, the curve and regions should extend to zero, but numerical error due to the removable singularity prevents this.
It is tempting to conjecture that this ansatz is the actual form of $m_{2}$ since this would agree with the critical behaviour, curve fitting evidence, and the GUE limit at $g$ goes to zero, and it bears similarity to the solution of similar models in [khalkhali2024coloured]. However, we can prove the solution is not of this form using the results of Section 4.1.2. Observe that
@@
The formal matrix model 8 has the asymptotic expansion of
@@ 4.2 When $(g,\alpha,\beta)=(g,-g,g)$
Moreover, we conjecture that the series expansion at zero of $m_{2}$ as a function of $g$ is the same as in the $(g,g,g)$ case, from which the above conjecture can be deduced. Our argument for this is as follows. Solving the SDE gives the same equations for the moments in terms of $m_{2}$, except for a subset of equations whose sign is flipped compared to the $(g,g,g)$ case; however, one can observe that the latter can only happen for those moments whose $g\rightarrow 0$ limit is zero. This is because both models converge to the GUE case when $g\to 0$.
@@
A natural corollary of this observation is that all our reasoning about the critical behavior of $m_{2}$ as a function of $g$ carries over from Section 4.1 into this case, and similarly, the critical exponent and asymptotic series expansion does so as well. We emphasize that there is no a priori reason to believe that the above conjecture is true.
@@ 4.3 When $(g,\alpha,\beta)=(g,g,-g)$
The search space of this model seems to be two, relying on $m_{2}$ and $m_{4}$. One can see in Figure 9 clear signs of a stable convergence to a solution for $g\geq 0$. The potential critical point in the regions corresponds to the peak of the surface in the figure. As the Hankel matrix size increases to sixteen by sixteen, the convergence slows.
Figure 7: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model for various sizes of the submatrix of the positive-definite Hankel matrix. The stripped region near zero is the result of numerical error. Note these plots are actually quite thin slices, which may be hard to discern from the presented fixed angle.
Ideally, we would hope for a finer curve; however, to deduce estimates of critical behaviour such a bootstrapped solution will suffice. Let us consider the level curves of the solution space for fixed $m_{4}$. Despite these level curves having different slopes, the power of the line of best fit remains approximately 0.85. For example, in Figure 8 the power is 0.886811 for the left subfigure when $m_{4}=3$ and 0.83865 for the right subfigure when $m_{4}=3$. Therefore even without ideal convergence, we can estimate there is a critical point for $m_{2}$ of this model near $g=0.165$ with a critical exponent of 0.85. This allows us to estimate the string susceptibility exponent since one can compute that
$\frac{d}{dg}F_{0}=-\frac{1}{2}m_{4}-\frac{1}{2}m_{ABAB}+m_{AABB}=\frac{m_{2}-1}{2g}.$ (12)
Hence, the estimated growth power of $\frac{d^{2}}{dg^{2}}F_{0}$ near the critical point is -0.15, making our estimate for the string susceptibility exponent $\gamma\approx 0.15$. This estimate is significantly distant from the usual two string susceptibility exponents to rule out either, and may represent a new continuum limit.
Figure 8: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model on the level curves $m_{4}=3$ on the left and $m_{4}=5$ on the right for a submatrix size twenty-two by twenty-two. The red curves are the lines of best fit.
@@ 4.4 When $(g,\alpha,\beta)=(-g,g,g)$
This model is very similar to $(g,\alpha,\beta)=(g,g,-g)$ configuration. The search space of this model also appears to be two, relying on $m_{2}$ and $m_{4}$. The method’s convergence also slows after a Hankel submatrix size of sixteen by sixteen is used.
Figure 9: Bootstrapping results for the $(g,\alpha,\beta)=(-g,g,g)$ 2-matrix model for various sizes of the submatrix of the positive definite Hankel matrix. The stripped region near zero is the result of numerical error.
@@
Figure 10: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model on the level curves $m_{4}=3$ on the left and $m_{4}=5$ on the right for a submatrix size twenty-two by twenty-two. The red curves are the lines of best fit.
@@ 5 The 3-matrix model
which is a generalization of the three-colour model studied in [eynard1998iterative, kostov2002exact]. As far as the authors can tell, this model is unsolved in the literature. Its search space dimension seems to be two and all moments are generated through the SDE from $m_{1}$ and $m_{2}$.
@@
Figure 11: The bootstrapped estimate of the 3-matrix model (13) for a fourteen by fourteen submatrix of the Hankel matrix with a lexicographical basis.
@@
$-\frac{d}{dg}\ln Z=\frac{1}{3}\langle\operatorname{Tr}(A^{3}+B^{3}+C^{3})\rangle+\langle\operatorname{Tr}ABC+\operatorname{Tr}ACB\rangle.$
@@ 6 Conclusion and discussion
In this paper, we investigated the critical phenomena of several multi-matrix models obtaining estimates for critical points and critical exponents via bootstrapping with positivity. In particular, for various 2-matrix models we find evidence of a string susceptibility exponent of $\gamma=1/2$ pulacing those models potentially in the universality class of the Continuum Random Tree. For other 2-matrix models, we found evidence that $\gamma\approx 0.15$ and $0.05$, which are far the critical exponents of any model known to the authors, certainly not the most typical $\gamma=1/2$ or $-1/2$. Such models may have a new continuum limit. This critical behaviour is of great interest in various approaches to quantum gravity such as 2d conformal field theory, tensor models, and Dynamical Triangulations. It is the belief of the authors that the critical exponents of many unsolved matrix models can be accurately estimated by bootstrapping with positivity, the work done here serving as a first example. In theory, such bootstrapping techniques could also be applied to tensor models that can be transformed into matrix models [lionni2018colored], allowing for positivity constraints of the matrix models.
@@ A.2 The moments
The following relations were found by solving the SDE in Mathematica:
@@
$m_{AABB}$ $=\frac{1-m_{2}}{4g}$
$m_{ABAB}$ $=\frac{-4gm_{2}^{2}-m_{2}+1}{4g}$
@@
$m_{AAAABB}$ $=\frac{-4g^{2}m_{2}^{3}-4gm_{2}^{2}+8gm_{2}+m_{2}-1}{16g^{2}}$
$m_{AAABAB}$ $=\frac{-4g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
$m_{AABAAB}$ $=\frac{12g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
@@
$m_{AAAAAABB}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+12gm_{2}^{2}-24gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{AAAAABAB}$ $=\frac{16g^{3}m_{2}^{4}+32g^{2}m_{2}^{3}-28g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AAAABAAB}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{AAAABBBB}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+8gm_{2}^{2}-16gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AAABAAAB}$ $=\frac{-48g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-8gm_{2}^{2}+8gm_{2}-4g-m_{2}+1}{64g^{3}}$
$m_{AAABABBB}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{AAABBABB}$ $=\frac{16g^{3}m_{2}^{4}+4g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AABABABB}$ $=\frac{-16g^{3}m_{2}^{4}-28g^{2}m_{2}^{2}-4gm_{2}^{2}-8gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{AABABBAB}$ $=\frac{-16g^{3}m_{2}^{4}-32g^{2}m_{2}^{3}+4g^{2}m_{2}^{2}-8gm_{2}^{2}+4g-m_{2}+1}{64g^{3}}$
$m_{AABBAABB}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}+8gm_{2}^{2}-24gm_{2}+12g-m_{2}+1}{64g^{3}}$
$m_{ABABABAB}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}-44g^{2}m_{2}^{2}-8gm_{2}^{2}-8gm_{2}+12g-m_{2}+1}{64g^{3}}$

### v2

@@ 3 The quartic model
Figure 1: Bootstrapped solution of the quartic Hermitian matrix model for $g>0$. The colours correspond to the the size of the submatrix of the Hankel matrix as follows: Light blue is for 2 by 2, dark blue is for 3 by 3, green is for 4 by 4, and gold is for 5 by 5. The black curve is the analytic solution.
@@
Figure 2: Bootstrapped solutions of the quartic Hermitian matrix model for $g<0$. The top left yellow region was computed with submatrices of size 6 by 6, the top right was computed with submatrices of size 7 by 7, and the bottom region was computed with submatrices of size 10 by 10. The critical point of this model is $-\frac{1}{12}=-.0833\overline{3}$.
@@
The second derivative of the free energy of the matrix model can then be expanded as
@@ 4 The 2-matrix models
This model can clearly be considered as a convergent integral when the coupling constant $g,\alpha,\beta>0$, or as a formal matrix integral that enumerates coloured maps otherwise. As far as the authors are aware, there is no known technique that can be used to solve this model in its full generality, but it has been solved in special cases. When $(g,\alpha,\beta)=(g,0,0)$, this is two uncoupled quartic ensembles. When $(g,\alpha,\beta)=(g,\alpha,0)$, this becomes the symmetric $ABAB$ model solved via the method of characteristic expansion in [kazakov1999two]. When $(g,\alpha,\beta)=(0,0,\beta)$, this matrix integral is sometimes known as the Hoppe model and has been solved via a reduction to the KP equation or saddle point method [kazakov1999d, berenstein2009multi]. For the particular configuration $(g,\alpha,\beta)=(g,\alpha,\alpha)$, this model was bootstrapped in [kazakov2022analytic], where a phase diagram was constructed.
@@ 4.1 When $(g,\alpha,\beta)=(g,g,g)$
By symbolically solving the SDE of the model in Mathematica up to a cutoff, all the the mixed moments we examined could be expressed solely in terms of the coupling constant $g$ and $m_{2}$; see Appendix A.2 for some examples. This suggests that this model has a search space of one. Despite the uncertain nature of this claim, by using the solutions mentioned above, we are able to produce several analytical results and conjectures in addition to numerical bootstrap estimates.
@@ 4.1.1 Bootstrap bounds
The second moment $m_{2}$ of the matrix model 7 is such that
$-\frac{1}{4g}-\frac{1}{4}\sqrt{\frac{8g+1}{g^{2}}}\leq m_{2}\leq-\frac{1}{4g}+\frac{1}{4}\sqrt{\frac{8g+1}{g^{2}}},$
for $g\in[-1/8,0)$. Additionally,
$-\frac{1}{4}\sqrt{\frac{8g+1}{g^{2}}}-\frac{1}{4g}<m_{2}<0,$
@@
$0<m_{2}<\frac{1}{4}\sqrt{\frac{8g+1}{g^{2}}}-\frac{1}{4g}$
@@
$\det\begin{bmatrix}1&0&0&m_{2}&0&m_{2}\\ 0&m_{2}&0&0&0&0\\ 0&0&m_{2}&0&0&0\\ m_{2}&0&0&m_{4}&0&m_{2,2}\\ 0&0&0&0&m_{2,2}&0\\ m_{2}&0&0&m_{2,2}&0&m_{4}\end{bmatrix}\geq 0,$
@@
Figure 3: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (5) overlaid as follows: yellow is 5 by 5, cyan is 9 by 9, and purple is 21 by 21. The subfigure on the right is only the latter constraint.
@@ 4.1.2 Series expansion at zero
$-\frac{d}{dg}F_{0}=\frac{1}{2}m_{4}+\frac{1}{2}m_{1,1,1,1}+m_{2,2}.$
@@ 4.1.3 Critical behaviour
Figure 4: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model for $g<0$. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (5) as follows: yellow is 5 by 5, cyan is 9 by 9, and purple is 21 by 21.
Figure 5: Bootstrapped estimates of $m_{2}$ of the $(g,\alpha,\beta)=(g,g,g)$ 2-matrix model for $g<0$. Each region corresponds to the positivity constraints of various sizes of submatrices of the Hankel matrix (5) as follows: yellow is 5 by 5, cyan is 9 by 9, and purple is 21 by 21. This is compared with the estimates from the series expansion of the second moment in Section 4.1.2. Each coloured line from left to right represents successively more terms in expansion (8).
@@
Figure 6: The bootstrapped solution compared with equation (10), in red, for an estimate of $g_{c}=-0.0502729$. As in other plots, the curve and regions should extend to zero, but numerical error due to the removable singularity prevents this.
It is tempting to conjecture that this ansatz is the actual form of $m_{2}$ since this would agree with the critical behaviour, curve fitting evidence, and the GUE limit as $g$ goes to zero, and it bears similarity to the solution of similar models in [khalkhali2024coloured]. However, we can prove that the solution is not of this form using the results of Section 4.1.2. Observe that
@@
The formal matrix model 7 has the asymptotic expansion of
@@ 4.2 When $(g,\alpha,\beta)=(g,-g,g)$
Moreover, it would follow from the above conjecture that the series expansion at zero of $m_{2}$ for the $(g,-g,g)$ model as a function of $g$ is the same as in the $(g,g,g)$ model. Our argument for this is as follows. Solving the SDE gives the same equations for the moments in terms of $m_{2}$, except for a subset of equations whose sign is flipped compared to the $(g,g,g)$ case; however, one can observe that the latter can only happen for those moments whose $g\rightarrow 0$ limit is zero. This is because both models converge to the GUE case when $g\to 0$.
@@
A natural corollary of this observation is that all our reasoning about the critical behaviour of $m_{2}$ as a function of $g$ carries over from Section 4.1 into this case, and similarly, the critical exponent and asymptotic series expansion does so as well. We emphasize that there is no a priori reason to believe that the above conjecture is true.
@@ 4.3 When $(g,\alpha,\beta)=(g,g,-g)$
The search space of this model seems to be two, relying on $m_{2}$ and $m_{4}$. One can see in Figure 9 clear signs of a stable convergence to a solution for $g\geq 0$. The potential critical point in the given regions corresponds to the peak of the surface in the figure. As the Hankel matrix size increases to sixteen by sixteen, the convergence slows.
Figure 7: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model for a 22 by 22 submatrix of the positive semi-definite Hankel matrix. The stripped region near zero is the result of numerical error. Note this plot is actually a quite thin slice, which may be hard to discern from the presented fixed angle.
Ideally, we would hope for a finer curve; however, to deduce estimates of critical behaviour such a bootstrapped solution will suffice. Let us consider the level curves of the solution space for fixed $m_{4}$. Despite these level curves having different slopes, the power of the line of best fit remains approximately 0.85. For example, in Figure 8 the power is 0.886811 for the left subfigure when $m_{4}=3$ and 0.83865 for the right subfigure when $m_{4}=3$. We can estimate there is a critical point for $m_{2}$ of this model near $g=0.165$ with a critical exponent of 0.85. This allows us to estimate the string susceptibility exponent since one can compute that
$\frac{d}{dg}F_{0}=-\frac{1}{2}m_{4}-\frac{1}{2}m_{1,1,1,1}+m_{2,2}=\frac{m_{2}-1}{2g}.$ (11)
Hence, the estimated growth power of $\frac{d^{2}}{dg^{2}}F_{0}$ near the critical point is -0.15, making our estimate for the string susceptibility exponent $\gamma\approx 0.15$. This estimate is significantly distant from the usual two string susceptibility exponents, and may represent a new continuum limit.
Figure 8: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model on the level curves $m_{4}=3$ on the left and $m_{4}=5$ on the right for a submatrix size 22 by 22. The red curves are the lines of best fit.
@@ 4.4 When $(g,\alpha,\beta)=(-g,g,g)$
This model is very similar to $(g,\alpha,\beta)=(g,g,-g)$ configuration. The search space of this model also appears to be two, relying on $m_{2}$ and $m_{4}$. The bootstrap convergence also slows after a Hankel submatrix size of sixteen by sixteen is used.
Figure 9: Bootstrapping results for the $(g,\alpha,\beta)=(-g,g,g)$ 2-matrix model for various sizes of the submatrix of the positive semi-definite Hankel matrix. The stripped region near zero is the result of numerical error. Note this plot is actually a quite thin slice, which may be hard to discern from the presented fixed angle.
@@
Figure 10: Bootstrapping results for the $(g,\alpha,\beta)=(g,g,-g)$ 2-matrix model on the level curves $m_{4}=3$ on the left and $m_{4}=5$ on the right for a submatrix size 22 by 22. The red curves are the lines of best fit.
@@ 5 The 3-matrix model
which is a generalization of the three-colour model studied in [eynard1998iterative, kostov2002exact]. As far as the authors can tell, this model is unsolved in the literature. For any word in $A,B$ and $C$ we denote the associated mixed moment in the large $N$ limit as
$m_{W}:=\lim_{N\rightarrow\infty}\frac{1}{N}\langle\operatorname{Tr}W\rangle.$
The search space dimension of this model seems to be two and all moments used can generated through the SDE from $m_{A}:=m_{1}$ and $m_{AA}:=m_{2}$.
@@
Figure 11: The bootstrapped estimate of the 3-matrix model (12) for a 14 by 14 submatrix of the Hankel matrix with a lexicographical basis.
@@
$-\frac{d}{dg}\ln\mathcal{Z}=\frac{1}{3}\langle\operatorname{Tr}(A^{3}+B^{3}+C^{3})\rangle+\langle\operatorname{Tr}ABC+\operatorname{Tr}ACB\rangle.$
@@ 6 Conclusion and discussion
In this paper, we investigated the critical phenomena of several multi-matrix models obtaining estimates for critical points and critical exponents via bootstrapping with positivity. In particular, for various 2-matrix models we find evidence of a string susceptibility exponent of $\gamma=1/2$ placing those models potentially in the universality class of the Continuum Random Tree. For other 2-matrix models, we found evidence that $\gamma\approx 0.15$ and $0.05$, which are far from the critical exponents of any model known to the authors, certainly not the most typical $\gamma=1/2$ or $-1/2$. Such models may have a new continuum limit. This critical behaviour is of great interest in various approaches to quantum gravity such as 2d conformal field theory, tensor models, and Dynamical Triangulations. It is the belief of the authors that the critical exponents of many unsolved matrix models can be accurately estimated by bootstrapping with positivity, the work done here serving as a first example. In theory, such bootstrapping techniques could also be applied to tensor models that can be transformed into matrix models [lionni2018colored], allowing for positivity constraints of the tensor models.
@@ A.2 The moments
The following relations were found by solving some large set of SDE in Mathematica:
@@
$m_{2,2}$ $=\frac{1-m_{2}}{4g}$
$m_{1,1,1,1}$ $=\frac{-4gm_{2}^{2}-m_{2}+1}{4g}$
@@
$m_{4,2}$ $=\frac{-4g^{2}m_{2}^{3}-4gm_{2}^{2}+8gm_{2}+m_{2}-1}{16g^{2}}$
$m_{3,1,1,1}$ $=\frac{-4g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
$m_{2,1,2,1}$ $=\frac{12g^{2}m_{2}^{3}+4gm_{2}^{2}+m_{2}-1}{16g^{2}}$
@@
$m_{6,2}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+12gm_{2}^{2}-24gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{5,1,1,1}$ $=\frac{16g^{3}m_{2}^{4}+32g^{2}m_{2}^{3}-28g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{4,1,2,1}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{4,4}$ $=\frac{-16g^{3}m_{2}^{4}+36g^{2}m_{2}^{2}+8gm_{2}^{2}-16gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{3,1,3,1}$ $=\frac{-48g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}+20g^{2}m_{2}^{2}-8gm_{2}^{2}+8gm_{2}-4g-m_{2}+1}{64g^{3}}$
$m_{3,1,1,3}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}-4gm_{2}^{2}-m_{2}+1}{64g^{3}}$
$m_{3,2,1,2}$ $=\frac{16g^{3}m_{2}^{4}+4g^{2}m_{2}^{2}-8gm_{2}+4g-m_{2}+1}{64g^{3}}$
$m_{2,1,1,1,1,2}$ $=\frac{-16g^{3}m_{2}^{4}-28g^{2}m_{2}^{2}-4gm_{2}^{2}-8gm_{2}+8g-m_{2}+1}{64g^{3}}$
$m_{2,1,1,2,1,1}$ $=\frac{-16g^{3}m_{2}^{4}-32g^{2}m_{2}^{3}+4g^{2}m_{2}^{2}-8gm_{2}^{2}+4g-m_{2}+1}{64g^{3}}$
$m_{2,2,2,2}$ $=\frac{16g^{3}m_{2}^{4}+16g^{2}m_{2}^{3}-12g^{2}m_{2}^{2}+8gm_{2}^{2}-24gm_{2}+12g-m_{2}+1}{64g^{3}}$
$m_{1,1,1,1,1,1,1,1}$ $=\frac{16g^{3}m_{2}^{4}-16g^{2}m_{2}^{3}-44g^{2}m_{2}^{2}-8gm_{2}^{2}-8gm_{2}+12g-m_{2}+1}{64g^{3}}$